# Coordinating Search-Informed Reasoning and Reasoning-Guided Search in Claim Verification

### v1

@@ 4.2 Baselines
We evaluate our method against the following baselines:
@@
A commonly used paradigm (Hu et al., 2025b; Jiang et al., 2024; Kamoi et al., 2023) involving three steps: decomposing a claim into sub-claims, verifying each independently, and aggregating the results. We utilize the decomposition module from Kamoi et al. (2023) and prompt the LLM for verification, aggregating final results with logical AND.
@@
Wang and Shu (2023) translates claims into First-Order Logic (FOL) clauses and applies FOL-guided reasoning over knowledge-grounded question-answer (QA) pairs. The QA pairs are grounded via an external API [^3] We re-implemented FOLK using the released repository and the same Google Search API (https://serpapi.com/) for knowledge grounding..
@@
4.3 Experimental Setup
We conduct all training experiments using the Qwen3-4B model (Qwen, 2025). All baseline methods, except for Search-o1, utilize GPT-4o (OpenAI, 2024b) as the underlying LLM. Search-o1 employs the QwQ-32B-preview model (Qwen, 2024). All methods operate in the Open-Book setting (Pan et al., 2023b), where no ground-truth evidence is provided beforehand, requiring each method to retrieve supporting evidence using top-$k$ ($k=3$) retrieval. We utilize the Wikipedia corpus provided by FlashRAG (Jin et al., 2024) for this purpose, indexed using a E5-small model [^4] We use the ‘intfloat/multilingual-e5-small’ model (Wang et al., 2024), which offers a favorable balance between memory efficiency and performance on the MTEB benchmark (Muennighoff et al., 2023). for dense retrieval.
For training, we set the rollout temperature to 1.0, the learning rate to $1\times 10^{-6}$, and the KL loss coefficient to 0.001. For more experimental details, please refer to Appendix A.5.
@@ 5.1 Main Result
As shown in Tables 1, our proposed method HARIS consistently outperforms all baseline methods across both the EX-FEVER and HOVER datasets, demonstrating superior multi-hop reasoning and evidence searching capabilities. Notably, HARIS achieves the highest F1 and accuracy scores across different hop counts, with particularly strong performance in the more challenging 3-hop and 4-hop settings. For example, on the HOVER dataset, HARIS achieves 62.80% accuracy in the 3-hop setting and 61.00% in the 4-hop setting, surpassing other strong GPT-4o-powered baselines.
@@ 5.2 Ablations
Multi-agent RL has shown strong performance in complex reasoning tasks (Wan et al., 2025), and reducing retrieval noise during intermediate steps is also known to benefit RAG systems like Search-o1 (Li et al., 2025). To assess the impact of our bi-level design, we compare HARIS with an RL-trained single-agent . As shown in Table 2, HARIS outperforms the single-agent setup across multiple datasets. This shows the advantage of decomposing reasoning and retrieval into specialized agents, each optimized for its specific role. The search agent efficiently provides relevant information to the reasoning agent, reducing noise and enhancing the verification capability.A qualitative case study in Appendix B.3 further illustrates this contrast on a shared example.
@@
We investigate the impact of coordinated agent training on model performance. Table 3 shows that increasing the number of coordination rounds ($N$) generally improves overall performance, indicating that this approach helps balance learning dynamics and maintain alignment between high-level and low-level agents. By allowing each agent to iteratively refine its abilities while maintaining consistency, coordinated training supports more effective collaboration over multiple training cycles.
@@
We also examine the impact of the reward metric used for training the search agent using 3,600 training examples over one epoch. Specifically, we compare the use of the conventional QA F1-score and the LLM-as-a-Judge approach as search agent’s rewards. We observed that search agent trained with the F1 reward tends to produce shorter responses, as the F1 metric encourages concise outputs closely aligned with the synthesized answers. In contrast, search agent trained with LLM-as-a-Judge is incentivized to generate more comprehensive, contextually relevant information. As shown in Figure 4 and Figure 5, using an F1-trained search agent leads the high-level reasoning agent to trigger more frequent searches—meaning the reasoning agent asks more follow-up questions. By comparison, the LLM-as-a-Judge approach results in fewer search agent calls but higher verification reward, reflecting that returning more thorough information better supports the overall verification reasoning process.
@@ 6 Case Study
As illustrated in Figure 6, we present an example of how HARIS resolves a complex claim. The displayed rollouts clearly shows the step-by-step search-informed reasoning and reasoning-driven search. The reasoning agent rollouts demonstrate systematic probing through multi-step reasoning, verifying plausible related actors. Meanwhile, the search agent rollouts showcase the ability to continuously refine queries, steering the search direction from Feodor Chin to Blizzard to gather sufficient information. In another reasoning agent rollout shown in Figure 7, the reasoning agent initially struggles to identify the correct Baby film, but the search agent’s response to the question "Who is David Huynh…" not only addresses the immediate question but also reveals the correct film context, guiding the reasoning agent toward the correct verification path. These examples highlight the collaborative nature of HARIS, where the reasoning agent continuously refines its understanding based on dynamically retrieved information until all critical connections are uncovered.
@@ 7 Conclusion
We propose Hierarchical Agent Reasoning and Information Search (HARIS), explicitly modeling the coordinated process of reasoning-driven searching and search-informed reasoning. By integrating specialized high-level reasoning and low-level search agents, HARIS effectively captures complex reasoning chains while reducing noise in evidence retrieval. Our approach demonstrates strong performance across challenging benchmarks, highlighting its effectiveness for comprehensive claim verification.
@@
While HARIS demonstrates strong performance, it has several limitations. Due to limited computational resources and high GPU requirements, we train only on a 4B model, though larger models are likely to achieve stronger performance. Additionally, our study focuses on binary classification (‘Support’ or ‘Refute’), while some claim verification benchmarks also include classes like ‘Neutral’ or ‘Not Enough Info’, which we have not explored.
@@ A.5.2 Hardware & Hyperparameter
All experiments are conducted on a 4xH20 server and a cluster with 8xA100 compute nodes. Our implementation is based on the verl framework (Sheng et al., 2024). Key hyperparameters include a rollout group size of 5, tensor parallel size (tp) of 2, batchsize of 48, temperature of 1.0, learning rate of 1e-6, and a KL coefficient of 0.001.
We host the retrieval services on 1 GPU using FastAPI. For high-level agent training, we host the low-level agent endpoint using vLLM (Kwon et al., 2023), which takes 1 more GPU, and use the rest for training high-level agent. Due to GPU memory constraints, we limit the maximum context length to 8192 tokens.
A.6 Prompt Template
@@
Appendix B Examples
B.1 Synthesized Training Data
Table B.1 shows two examples of synthesized training data for the search agent. Example 2 leverages human-annotated explanations as evidence. The synthesis process uses a Chain-of-Thought (CoT) prompting format, with the rationale field capturing GPT-4o’s intermediate reasoning before producing the final answer.
@@ B.1 Synthesized Training Data
{ "input": "94 episodes were in the original run of the HBO show in which the voice actress of Miss Spider played the character Charlotte York Goldenblatt.", "evidence": "Miss Flora Spider is the title character in the bestselling childrens books by David Kirk. She originally appeared in cartoon format in the movie \"Miss Spiders Sunny Patch Kids\" (with Brooke Shields providing Miss Spiders voice), which spun off the \"Miss Spiders Sunny Patch Friends\" series (with Kristin Davis providing her voice this time around), and \"Miss Spiders Sunny Patch Friends - Harvest Time Hop and Fly\" video game...", "explanation": null, "label": true, "question": "Who is the voice actress for the character \"Miss Spider\"?", "qa_response": { "rationale": "To produce the answer. We start by identifying the character \"Miss Spider\" and the voice actress associated with her. According to the evidence, Miss Spider is a character in the childrens books by David Kirk and has appeared in various media formats. The evidence states that Brooke Shields provided Miss Spiders voice in the movie \"Miss Spiders Sunny Patch Kids,\" while Kristin Davis provided her voice in the \"Miss Spiders Sunny Patch Friends\" series. Since the question asks for the voice actress for the character \"Miss Spider,\" we need to determine which context is being referred to. However, given the context of the claim and the evidence provided, it is most relevant to identify Kristin Davis as the voice actress for the series, as she is also the actress who played Charlotte York Goldenblatt in \"Sex and the City.\" Therefore, the answer to the question is Kristin Davis.", "answer": "Kristin Davis" } }
@@
{ "input": "Journey is an American rock band formed in 1974 by former members of band such as an American rock band formed in 1966.", "evidence": null, "explanation": "Journey is formed in 1973 by former members of band such as Steve Miller Band, not 1974. The Steve Miller Band is an American rock band formed in 1966.", "label": false, "question": "What American rock band was formed in 1966 and had members who later joined Journey?", "qa_response": { "rationale": "To produce the answer. We start by identifying the American rock band formed in 1966. According to the evidence, the Steve Miller Band was formed in 1966. Next, we need to determine if any members from this band later joined Journey. The evidence states that Journey was formed by former members of bands such as the Steve Miller Band. Therefore, the Steve Miller Band is the band in question.", "answer": "The Steve Miller Band." } }
@@
B.2 Synthesized Training Data
Table B.1 shows two examples of synthesized training data for the search agent. Example 2 leverages human-annotated explanations as evidence. The synthesis process uses a Chain-of-Thought (CoT) prompting format, with the rationale field capturing GPT-4o’s intermediate reasoning before producing the final answer.
B.3 Single-Agent & Multi-Agent Cases
@@
Figure 8: Example of a trained single agent rollout. The agent issues several queries but fails to retrieve sufficient evidence. Without a clear separation of roles, it struggles to iteratively refine its search direction and prematurely converges on partial evidence, ultimately leading to an incorrect verification.
Figure 9: Example of HARIS’s reasoning agent rollout. The reasoning and search agents coordinate effectively: the reasoning agent identifies uncertain links and delegates targeted queries, while the search agent retrieves precise evidence, enabling correct verification.

### v2

@@ 4.2 Baselines
We include the following baselines:
@@
A commonly used paradigm (Kamoi et al., 2023) involving: decomposing a claim into sub-claims, verifying each independently, and aggregating the results. We utilize the decomposition module from Kamoi et al. (2023) and prompt the LLM for verification, aggregating final results with logical AND.
@@
Wang and Shu (2023) translates claims into First-Order Logic (FOL) clauses and applies FOL-guided reasoning over knowledge-grounded question-answer (QA) pairs. The QA pairs are grounded via an external API [^5] https://serpapi.com/.
BiDeV
Liu et al. (2025) propose two prompt-based LLM agents for defusing vagueness and redundancy: the former clarifies latent information, while the latter removes redundant evidence.
@@
F1 LLM-as-a-Judge
F1 Acc F1 Acc
[…]
$\text{HOVER}_{\text{2hops}}$ 65.94 66.40 67.49 67.60
$\text{HOVER}_{\text{3hops}}$ 56.42 57.80 58.97 59.60
$\text{HOVER}_{\text{4hops}}$ 52.38 55.60 57.42 58.80
Table 4: Final performance comparison between HARIS with F1-trained and LLM-as-a-Judge trained search agent.
4.3 Experimental Setup
[…]
@@ 5.1 Main Result
As shown in Tables 1, our proposed method HARIS consistently outperforms all baseline methods across both the EX-FEVER and HOVER datasets, demonstrating superior multi-hop reasoning and evidence searching capabilities. Notably, HARIS achieves the highest F1 and accuracy scores across different hop counts, with particularly strong performance in the more challenging 3-hop and 4-hop settings. For example, on the HOVER dataset, HARIS achieves 62.80% accuracy in the 3-hop setting and 61.00% in the 4-hop setting, surpassing other strong GPT-4o-powered baselines. For direct comparison with Qwen3-4B results, see Table 12.
@@ 5.2 Ablations
Multi-agent RL has shown strong performance in complex reasoning tasks (Wan et al., 2025), and reducing retrieval noise during intermediate steps is also known to benefit RAG systems like Search-o1 (Li et al., 2025). To assess the impact of our bi-level design, we compare HARIS with an RL-trained single-agent . As shown in Table 2, HARIS outperforms the single-agent setup across multiple datasets. This shows the advantage of decomposing reasoning and retrieval into specialized agents, each optimized for its specific role. The search agent efficiently provides relevant information to the reasoning agent, reducing noise and enhancing the verification capability. A qualitative case study in Appendix C.4 further illustrates this contrast on a shared example.
@@
We study how the number of coordination rounds ($N$) affects model performance. As shown in Table 3, increasing $N$ generally leads to stronger performance. It shows that this approach helps balance learning dynamics and maintain alignment between high-level and low-level agents. By allowing each agent to iteratively refine its abilities while maintaining consistency, coordinated training supports more effective collaboration over multiple training cycles.
@@
We examine the impact of the reward metric used for training the search agent using 3,600 training examples over one epoch. Specifically, we compare conventional QA F1-score and LLM-as-a-Judge as rewards. We find that F1-trained agents tend to generate more concise responses, as F1 favors answers closely matching the reference. In contrast, LLM-as-a-Judge rewards encourage more comprehensive and contextually relevant outputs. As shown in Figure 4 and  5, F1-based agents prompt the reasoning agent to trigger more searches, resulting in more follow-up questions. In comparison, LLM-as-a-Judge reduces search calls but achieves higher verification reward, indicating more thorough information improves the overall reasoning process. As shown in Table 4, the performance results suggest using LLM-as-a-Judge trained HARIS consistently improved the performance compared to F1-trained. On average it improves over 3% performance.
More experiments can be found in Appendix B.
@@ 6 Case Study
Figure 6 illustrates how HARIS resolves a complex claim through step-by-step, search-informed reasoning. The reasoning agent systematically probes plausible related actors, while the search agent continuously refines queries, shifting the search from Feodor Chin to Blizzard to gather sufficient evidence. In another example (Figure 7), the reasoning agent initially struggles to identify the correct Baby film, but the search agent’s response about "Who is David Huynh…" provides crucial context, steering reasoning toward the correct verification. These cases highlight HARIS’s collaborative process, with the reasoning agent refining its understanding as new information is retrieved until all critical connections are uncovered.
@@ 7 Conclusion
We propose Hierarchical Agent Reasoning and Information Search (HARIS), explicitly modeling the coordinated process of reasoning-driven searching and search-informed reasoning. By integrating high-level reasoning and low-level search agents, HARIS effectively captures complex reasoning chains while reducing noise in evidence retrieval. Our approach demonstrates strong performance across challenging benchmarks, highlighting its effectiveness for comprehensive claim verification.
@@
While HARIS demonstrates strong performance, due to limited computational resources, we train only on a 4B model. Using larger models are likely to achieve even stronger performance. Additionally, our study focuses on binary claim verification (‘Support’ or ‘Refute’). While some benchmarks include additional classes such as ‘Neutral’ or ‘Not Enough Info,’ we do not explore them here. Claim verification is a key area within fact-checking, we do not explore tasks such as open-domain QA or counterfactual detection, as these differ substantially from multi-hop claim verification. Notably, our binary setting is consistent with strong baselines such as ProgramFC, QACheck, and FOLK in claim verification.
@@ A.4 Search Agent Training Data Synthesis
HOVER EX-FEVER
2-hops 3-hops 4-hops 2-hops 3-hops
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
RAG 57.4757.4757.47 57.6057.6057.60 49.4949.4949.49 52.4052.4052.40 50.7450.7450.74 53.0053.0053.00 68.6768.6768.67 68.8068.8068.80 68.5468.5468.54 68.6068.6068.60
Decomp-Verify 59.6759.6759.67 60.8060.8060.80 47.1347.1347.13 53.7153.7153.71 43.5943.5943.59 52.8052.8052.80 63.0963.0963.09 64.6064.6064.60 54.7654.7654.76 59.0059.0059.00
ProgramFC 55.4155.4155.41 55.6055.6055.60 50.3950.3950.39 51.2051.2051.20 51.1551.1551.15 52.8052.8052.80 56.6056.6056.60 56.8056.8056.80 56.3156.3156.31 57.6057.6057.60
BiDeV 59.3859.3859.38 62.4062.4062.40 50.5550.5550.55 56.2056.2056.20 42.0942.0942.09 52.6052.6052.60 58.7358.7358.73 62.2062.2062.20 51.6051.6051.60 57.6057.6057.60
QACheck 55.9055.9055.90 56.2056.2056.20 48.0848.0848.08 50.2050.2050.20 51.0751.0751.07 52.8052.8052.80 55.8455.8455.84 56.2056.2056.20 60.1460.1460.14 60.4060.4060.40
FOLK 61.0861.0861.08 61.4061.4061.40 59.0759.0759.07 59.2059.2059.20 57.6557.6557.65 58.2058.2058.20 80.1280.1280.12 80.2080.2080.20 73.9373.9373.93 74.2074.2074.20
HARIS 69.31 69.40 62.33 62.80 59.84 61.00 80.12 80.20 73.93 74.20
Table 5: Direct performance comparison of different methods (Qwen3-4B based) with HARIS. [^12] For Search-o1, its official implementation is specifically designed for QwQ reasoning models and is not directly configurable with Qwen3 models.
@@ A.5.2 Hardware & Hyperparameter
All experiments, including HARIS and the baselines, were conducted on a server with 4×H20 GPUs and a cluster of 8×A100 nodes. Our implementation is based on the verl framework (Sheng et al., 2024). Key hyperparameters include: rollout group size of 5, tensor parallel size (tp) of 2, batch size of 48, temperature of 1.0, learning rate of 1e-6, and KL coefficient of 0.001.
HARIS experiments were mainly run on the 4×H20 server. Retrieval services were hosted on a single GPU using FastAPI. For reasoning agent training, we used vLLM (Kwon et al., 2023) to serve the search agent endpoint on one GPU, while the remaining GPUs were allocated to high-level agent training. Two GPUs for training the reasoning agent and one GPU for the search agent inference service. In the single-agent setting, two GPUs were used for training. Due to GPU memory constraints, we set the maximum context length to 8192 tokens.
Appendix B Additional Experiments
B.1 Direct comparison
To enable direct comparison on the same base model, we run the baselines using Qwen3-4B as the base LLM. The results are summarized in Table 12. As shown, when using the same base LLM, HARIS significantly outperforms the baselines, demonstrating its effectiveness.
B.2 Supervised Finetuning
We provide an experiment comparing HARIS with supervised finetuning (SFT). Specifically, in an explainable fact-checking setting, we enabled the thinking mode of Qwen3-4B to generate responses using the same training set as HARIS. To ensure the model learns the correct target sequence, for each claim in the training data, we repeatedly sampled responses until the final prediction was correct. The prompt used can be found in Table 9. Training epochs and learning rates were kept the same as HARIS’s setting. The performance results are summarized in Table 6. Overall, HARIS outperforms supervised fine-tuning across all datasets and hop settings.
B.3 ReAct & Model Scaling
One might be concerned that decoupling the search and reasoning agents primarily compensates for the limitations of smaller models (such as our 4B backbone). However, in our main experiments, the Search-o1 baseline employs a larger 32B model, yet it still underperforms compared to HARIS. To further investigate the effect of model scaling, we implemented a ReAct LLM Agent baseline and conducted experiments using Qwen3-4B, 8B, and 14B. In this setup, the agent performs Wikipedia searches and leverages the retrieved documents as observations. The F1 results are presented in Table 7. As shown, simply increasing the model size does not always result in substantial performance gains. These results suggest that our multi-agent, decoupled approach offers distinct advantages.
SFT HARIS
F1 Acc F1 Acc
[…]
$\text{HOVER}_{\text{2hops}}$ 57.05 59.40 69.31 69.40
$\text{HOVER}_{\text{3hops}}$ 42.03 51.40 62.33 62.80
$\text{HOVER}_{\text{4hops}}$ 41.24 52.00 59.84 61.00
Table 6: Performance comparison between supervised finetuning(SFT) and HARIS.
ReAct HARIS
4B 8B 14B 4B
[…]
$\text{HOVER}_{\text{2hops}}$ 63.59 63.19 64.57 69.31
$\text{HOVER}_{\text{3hops}}$ 54.94 58.19 55.76 62.33
$\text{HOVER}_{\text{4hops}}$ 51.24 53.17 56.02 59.84
Table 7: F1 performance comparison between HARIS and ReAct agent based on different model sizes.
Appendix C Prompts & Examples
C.1 Prompt Template
@@
Prompt Template for Single Agent
You are a helpful assistant tasked with verifying the truthfulness of a claim step by step, with the help of the wikipedia search tool. Given a claim, you need to first think about the reasoning process in the mind and then provide the boolean verification result (Support or Refute). During thinking, you can invoke the wikipedia search tool to search for fact information about specific topics if needed. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags respectively, and the search query and result are enclosed within <search> </search> and <result> </result> tags respectively…
Prompt Template for Supervised Finetuning
<|im_start|>user
Given a claim and its retrieved evidence, determine whether the claim is ’Support’ or ’Refute’.
Claim: claim
Evidence: retrieved_evidence
Wrap your final answer in <answer> and </answer> tags (e.g. <answer>Support</answer> or <answer>Refute</answer>)<|im_end|>
<|im_start|>assistant
<think>
Table 9: Prompt templates for the single agent and supervised finetuning.
@@
C.2 Synthesized Training Data
Table C.2 shows two examples of synthesized training data for the search agent. Example 2 leverages human-annotated explanations as evidence. The synthesis process uses a Chain-of-Thought (CoT) prompting format, with the rationale field capturing GPT-4o’s intermediate reasoning before producing the final answer.
@@ C.2 Synthesized Training Data
C.3 Synthesized Training Data
Table C.2 shows two examples of synthesized training data for the search agent. Example 2 leverages human-annotated explanations as evidence. The synthesis process uses a Chain-of-Thought (CoT) prompting format, with the rationale field capturing GPT-4o’s intermediate reasoning before producing the final answer.
C.4 Single-Agent & Multi-Agent Cases
@@
Figure 8: Example of a trained single agent rollout. The input claim ‘Sam Stockman stars in a series that depicted a killer. The killer was copying the actions of Jack the Ripper.’ is correct. The agent issues several queries but fails to retrieve sufficient evidence. Without a clear separation of roles, it struggles to iteratively refine its search direction and prematurely converges on partial evidence, ultimately leading to an incorrect verification.
Figure 9: Example of HARIS’s reasoning agent rollout. The input claim ‘Sam Stockman stars in a series that depicted a killer. The killer was copying the actions of Jack the Ripper.’ is correct. The reasoning and search agents coordinate effectively: the reasoning agent identifies uncertain links and delegates targeted queries, while the search agent retrieves precise evidence, enabling correct verification.
Appendix D Human Evaluation Details
As described in Section 3.3.3, we conducted a human evaluation to assess the reliability of our LLM-as-a-Judge setup. We sampled 150 questions from a held-out set of synthesized QA data, using the trained HARIS search agent to gather information for each. Two annotators from the Prolific platform [^13] https://www.prolific.com/, each paid £20, independently evaluated 75 responses following the same guidelines as the LLM-as-a-Judge. They judged whether the retrieved information was sufficient and useful for deriving the correct answer. The results showed a Cohen’s Kappa of 0.81 and a 93.3% agreement rate, indicating strong consistency. These findings confirm that our LLM-as-a-Judge metric closely aligns with human judgments. An example of the annotation panel is shown in Figure 10. For data consent, we selected the AI task annotation category on the platform, and annotators were informed that the collected data would be used to evaluate LLM outputs.
Figure 10: Preview of the human evaluation panel.